# MIB: A Mechanistic Interpretability Benchmark

### v1

@@ E.5.4 Indirect Object Indentification
After an interchange intervention is performed on all four heads with one of these counterfactuals, Wang et al. (2023) consider the token/position to have value 1 if unchanged and value -1 if inverted. [^13] They also included a value of 00 for the token signal of a random new token, but we leave out this condition for our experiments. Using these binary signals as inputs and the logit difference between indirect object and subject as outputs, they perform a regression and find that $2.31\texttt{PositionSignal}+0.99\texttt{TokenSignal}$ is the best predictive model of logit difference. On our dataset, we perform the same intervention experiments and fit a linear model, finding that $0.069+2.018\texttt{PositionSignal}+0.687\texttt{TokenSignal}$ is the best predictor of logit difference.
@@
Without an intervention, the $S_{\text{Tok}}$ and $S_{\text{Pos}}$ variables will always match the text input, and the logit diff is predicted to be $0.069+2.018+0.687$. When only $S_{\text{Tok}}$ is intervened on, the logit diff is predicted to be $0.069+2.018-0.687$. When only $S_{\text{Pos}}$ is intervened on, the logit diff is predicted to be $0.069-2.018+0.687$. If both are intervened on, then the logit diff is predicted to be $0.069-2.018-0.687$
@@
We conduct a brute-force search by aligning each of the two causal variables with every possible subset of the four S-inhibition heads. In Table 17) we report the mean-squared error for the alignments. In the main text, we report the alignment of $S_{\text{Pos}}$ to heads 7.3, 7.9, and 8.6 and $S_{\text{Tok}}$ to the head 8.10.
LogitDifferenceSubjectTokenSubjectPositionTextPrompt×0.687×2.018+0.069
(a) General causal model for indirect object identification
LogitDifference = 3.16SubjectToken = 1SubjectPosition = 1 TextPrompt: ”As Carl and Maria left the consulate, Carl gave a fridge to” ×0.687×2.018+0.295
(b) Specific example showing prediction for ”Carl”
Figure 18: Causal model for indirect object identification. The TextPrompt is processed to extract the subject token $S_{\text{Tok}}$ and the subject position $S_{\text{Pos}}$. The output variable mechanism (1) compares the token and position to the input, and determines whether the token and position were inverted or the same and sets PositionSignal and TokenSignal to 1 and -1 accordingly and (2) computes $0.069+0.687\texttt{TokenSignal}+2.018\texttt{PositionSignal}$, with position having a stronger influence than token identity.
@@
((7, 3), (7, 9), (8, 6), (8, 10)) 3.82 13.4
((7, 3), (7, 9), (8, 6)) 2.25 6.34
((7, 3), (7, 9), (8, 10)) 2.96 6.2
((7, 3), (7, 9)) 4.96 3.07
((7, 3), (8, 6), (8, 10)) 2.38 7.37
((7, 3), (8, 6)) 3.83 4.44
((7, 3), (8, 10)) 4.68 3.26
((7, 3)) 11.08 3.83
((7, 9), (8, 6), (8, 10)) 2.73 9.42
((7, 9), (8, 6)) 2.83 4.7
((7, 9), (8, 10)) 3.43 4.0
((7, 9)) 7.2 2.62
((8, 6), (8, 10)) 2.77 5.29
((8, 6)) 5.28 3.92
((8, 10)) 6.78 2.67
Appendix F Table of Notation
Table 18: The notation used throughout the paper, grouped by track.
Track Symbol Meaning
Shared $\mathcal{N}$ The full computation graph; a neural network
Circuit Localization $\mathcal{C}$ A circuit $\in\mathcal{N}$
$u$ A node in the computation graph. Could be an MLP or an attention head
$(u,v)$ An edge from node $u$ to node $v$
$N_{u}$ The number of neurons in node $u$, usually equal to $d_{\text{model}}$
$N_{\mathcal{C}}$ The set of all neurons in $\mathcal{C}$
Causal Variable Localization $\mathcal{H}$ A high-level causal model
$X$ A variable in $\mathcal{H}$
$b$ An base input to the neural network
$\mathcal{D}$ A dataset containing $(b,c)$ pairs

### v2

@@ F.5.4 Indirect Object Indentification
After an interchange intervention is performed on all four heads with one of these counterfactuals, Wang et al. (2023) consider the token/position to have value 1 if unchanged and value -1 if inverted. [^13] They also included a value of 00 for the token signal of a random new token, but we leave out this condition for our experiments. Using these binary signals as inputs and the logit difference between indirect object and subject as outputs, they perform a regression and find that $2.31\texttt{PositionSignal}+0.99\texttt{TokenSignal}$ is the best predictive model of logit difference. On our dataset, we perform the same intervention experiments and fit a linear model, finding that $0.048+2.005\texttt{PositionSignal}+0.768\texttt{TokenSignal}$ is the best predictor of logit difference.
@@
Without an intervention, the $S_{\text{Tok}}$ and $S_{\text{Pos}}$ variables will always match the text input, and the logit diff is predicted to be $0.048+2.005+0.768$. When only $S_{\text{Tok}}$ is intervened on, the logit diff is predicted to be $0.048+2.005-0.768$. When only $S_{\text{Pos}}$ is intervened on, the logit diff is predicted to be $0.048-2.005+0.768$. If both are intervened on, then the logit diff is predicted to be $0.048-2.005-0.768$
@@
We conduct a brute-force search by aligning each of the two causal variables with every possible subset of the four S-inhibition heads. In Table 18 we report the mean-squared error for the alignments. In the main text, we report the alignment of $S_{\text{Pos}}$ to heads 7.3, 7.9, and 8.6 and $S_{\text{Tok}}$ to the head 8.10.
@@
((7, 3), (7, 9), (8, 6)) 2.45 6.83
((7, 3), (7, 9), (8, 10)) 3.08 6.79
((7, 3), (7, 9)) 5.42 3.28
((7, 3), (8, 6), (8, 10)) 2.52 8.08
((7, 3), (8, 6)) 4.17 4.82
((7, 3), (8, 10)) 4.92 3.51
((7, 3)) 12.03 4.07
((7, 9), (8, 6), (8, 10)) 2.90 10.20
((7, 9), (8, 6)) 3.15 5.05
((7, 9), (8, 10)) 3.63 4.33
((7, 9)) 7.90 2.80
((8, 6), (8, 10)) 3.00 5.71
((8, 6)) 5.81 4.21
((8, 10)) 7.28 2.82